OpenStudy (anonymous):

how do i do this i will give a medal The population of a town has been growing, following the equation P = 400 t + 5500, where t is years after 2010. The number of restaurants in the town has been growing according to the equation R = 2 t + 25. Complete an equation for the number of restaurants per capita (per person) Restaurants per capita:______

OpenStudy (danjs):

R / P

OpenStudy (anonymous):

This was part 2 but i only need to answer one of the question How many restaurants per capita does this model predict for the year 2017?

OpenStudy (anonymous):

anyone have any idea

OpenStudy (wolf1728):

39 Restaurants per 8,300 people as of 2017 Found just by trial and error using spreadsheet. Guess you'll like a formula though huh?

OpenStudy (anonymous):

only if u want to post it but thanks

OpenStudy (wolf1728):

I'll have to create it - might take a while.

OpenStudy (anonymous):

it is still saying that answer is wrong

OpenStudy (wolf1728):

population = 400*(Year -2010) + 5500 restaurants = 2 * (Year -2010) + 25 My answer is wrong? If they want restaurants per capita then I guess to be technical the answer is 1 restaurant per 212.8205128205 people (that's 8,300 divided by 39)

OpenStudy (wolf1728):

Restaurants per capita = [400*(Year -2010) + 5500] / [2 * (Year -2010) + 25]

OpenStudy (wolf1728):

So for 2017: Restaurants per capita = [400*(7) + 5500] / [2 * (7) + 25] Restaurants per capita = [2800 + 5500] / [14 + 25] Restaurants per capita = [2800 + 5500] / [14 + 25] Restaurants per capita = (8,300 / 39) = 212.8205128205

OpenStudy (wolf1728):

Matthew, was I right?